consider a linear temperature scale x have -75.5º X of boiling point of water and -190ºX freezing point. what is the temperature of 2.00ºC on the X scale.
1
Expert's answer
2020-08-24T11:55:50-0400

Let's consider X as a new temperature in linear scale and t - real temperature. Since the scale is linear the only possible connection between X and t is "X = at+b", where a and b are constants. In is enough to know 2 points to completely define this equation. We know that freezing point for water is "t = 0^\circ C" and "X = -190^\circ". Putting this into the equation, we have

"-190 = a \cdot 0 + b \Rightarrow b= - 190"

Also, we know that for boiling point "t=100^\circ C, X = -75.5^\circ."

"-75.5 = a \cdot 100 - 190 \Rightarrow a = 1.145"

So, "X(t) = 1.145t - 190".

"X(2.0) = 1.145 \cdot 2 - 190 = -187.71^\circ".

Answer: "-187.71^\circ" .